# DynamicID: Zero-Shot Multi-ID Image Personalization with Flexible Facial Editability

### v1

@@ 4.2 Comparation Results
Multi-ID personalization. To validate the efficacy of our DynamicID for multi-ID personalization, we conducted comprehensive comparisons with existing state-of-the-art methods, including those proposed in [77, 72, 24]. Figure 2 showcases a visual comparison of qualitative results, while Table 2 provides the quantitative analysis. These evaluations underscore that our method uniquely enables independent, fine-grained editing of individual facial features. In contrast, FastComposer, Face-diffuser, and Uniportrait face challenges with direct facial replication. Furthermore, while maintaining comparable prompt consistency with baseline approaches, our method significantly enhances the visual quality of generated images. The higher score of Uniportrait in the FaceSim metric compared to our method stems from the same reason as discussed in the context of single-ID personalization. More visual examples for three-ID images are shown in Fig. 10.
@@ 4.3 Ablation study
Motivation Verification: Fig. 7 demonstrates that during the entire denoising process, only the queries corresponding to facial regions exhibit strong activation to incorporate reference facial features. In contrast, queries corresponding to background regions remain non-activated, thereby preserving the behavior of the T2I model in these regions without interference.
@@
Figure 8: Ablation study on SAA. (a) The SAA preserves the original model’s behavior when queries are adaptively activated. (b) Additionally, it allows for the manual enhancement of query activation levels in specific regions, which can be specified by the user using bounding boxes, thereby enabling zero-shot layout control and multi-ID personalized generation.
@@ 5 Conclusion
We present DynamicID, a novel framework that enables high identity fidelity and flexible facial editing in personalized image generation across single- and multi-ID scenarios. Our approach introduces two key components: 1) Semantic-Activated Attention (SAA), which enhances feature integration through a query-level activation gating mechanism, thereby minimizing disruption to the original model and enabling zero-shot multi-ID personalization; 2) Identity-Motion Reconfigurator (IMR) that achieves superior facial editability via disentanglement and re-entanglement of identity and motion features in facial space. The framework employs task-decoupled training to eliminate dependency on specialized datasets, complemented by our newly developed VariFace-10k dataset containing diverse facial variations across 10k identities. Extensive experiments demonstrate state-of-the-art performance, validating the method’s effectiveness and broad application potential in personalized generation.

### v2

@@ 4.1 Experimental Setup
Figure 6: Visual comparison on Multi-ID Personalization. Current methods demonstrate constraints in multi-ID generation and unsatisfying editability. In contrast, our method enables flexible facial editability and high image quality in such scenarios.
@@ 4.2 Comparation Results
Multi-ID personalization. To validate the efficacy of our DynamicID for multi-ID personalization, we conducted comprehensive comparisons with existing state-of-the-art methods, including those proposed in [78, 73, 24, 34, 83]. Figure 6 showcases a visual comparison of qualitative results, while Table 2 provides the quantitative analysis. These evaluations underscore that our method uniquely enables independent, fine-grained editing of facial features. In contrast, FastComposer, Face-diffuser, Uniportrait, OMG-InstantID and StoryMaker face challenges with direct facial replication. Furthermore, while maintaining comparable prompt consistency with baseline approaches, our method significantly enhances the visual quality of generated images. The higher score of Uniportrait and OMG-InstantID in the FaceSim metric compared to our method stems from the same reason as discussed in the context of single-ID personalization. More visual examples for three-ID images are shown in Fig. 10.
@@ 4.3 Ablation study
Motivation Verification: Fig. 7 shows that during the entire denoising process, only the queries corresponding to facial regions exhibit strong activation. In contrast, queries corresponding to background regions remain non-activated, thereby preserving the behavior of the T2I model in these regions without interference.
@@
Figure 8: Ablation study on SAA. (a) The SAA preserves the original model’s behavior when queries are adaptively activated. (b) Moreover, it allows for the manual enhancement of query activation levels within user-defined spatial regions via bounding box annotations, thereby enabling zero-shot layout control and multi-ID personalized generation.
@@
5 Limitations
Although our plug-and-play framework fundamentally supports all text-to-image models, this work primarily uses Stable Diffusion v1.5 [55], which limits performance when generating faces near image edges [69] or handling complex prompts. Future work will explore scaling via larger datasets, broader model adaptation, and version releases.
@@ 6 Conclusion
We present DynamicID, a novel framework that enables high identity fidelity and flexible facial editing in personalized image generation across single- and multi-ID scenarios. Our approach introduces two novel components: SAA and IMR and employs task-decoupled training to eliminate dependency on specialized datasets, complemented by our newly developed VariFace-10k dataset containing diverse facial variations across 10k identities. Extensive experiments demonstrate state-of-the-art performance, validating the method’s effectiveness and broad application potential in personalized generation.